OpenStudy (anonymous):

HELP ME!! Evaluate: -4-4x(-6)

OpenStudy (anonymous):

simplify step-by-step. 4-4x(-6) =-4 +24x answer=?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

actually 24x-4

OpenStudy (anonymous):

can i ask you like 2 more?? @A.K.C.15 i really need some help i have to get this done

OpenStudy (anonymous):

yes and can you find somebody to help with mine on my page

OpenStudy (anonymous):

\[5-(-15)\div(-5)\]

OpenStudy (anonymous):

i dont know how to do really anything man i just signed up today all ik how to do is ask things /.\

OpenStudy (anonymous):

2 is the answer

OpenStudy (anonymous):

i tried posting it on that comment thing it got ignored :/ ok thanks! one more plz?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-4x3-(-6)

OpenStudy (anonymous):

-4x3+6

OpenStudy (anonymous):

do i multiply the -4 and 3 first?

OpenStudy (anonymous):

thats -12...then id add 6...and that would beeee -6???

OpenStudy (anonymous):

i dont get it...